Find the center and radius of the given circle and graph it.

2a. x^2+(y+5)^2=4

2b. x^2-6x+y^2+14y+33=0
1
Expert's answer
2013-07-30T08:51:10-0400
2a. Center at (0,-5). Radius is 2

2b. x^2-6x+y^2+14y+33=0

x^2 - 6x +36 - 36 + y^2+14y+49 - 49 + 33 =0

(x-6)^2 + (y+7)^2 = 46

Hence it is a circle with radius of sqrt(46) and center at (6,7)